d^2u/dx^2 - d^2u/dy^2 = 2(e^(-x)) + 3sin2y


1
Expert's answer
2021-07-27T07:50:07-0400

"\frac{d^2u}{dx^2} - \frac{d^2u}{dy^2} = 2e^{-x} + 3\sin 2y"

This is an inhomogeneous linear PDE. Its general solution can be obtained as a sum of a partial solution and the general solution of the corresponding homogeneous linear PDE.


Consider the homogeneous linear PDE:

"\frac{d^2u}{dx^2} - \frac{d^2u}{dy^2} = 0"

This is a well-known wave equation, it has the general solution "u_{hom}= f_1(x+y) + f_2(x-y)" , where "f_1" and "f_2" are arbitrary functions.


Let's find a partial solution of the form "u_p=g(x)+h(y)" , where "g" , "h" are unknown functions.

"\frac{d^2u_p}{dx^2} - \frac{d^2u_p}{dy^2} =g''(x)-h''(y)=2e^{-x} + 3\sin 2y"

Hence

"g''(x)=2e^{-x}" and "g(x)=2e^{-x}" (partial solution)

"-h''(y)=3\sin 2y" and "h(y)=\frac{3}{4}\sin 2y" (partial solution)


Therefore, the general solution to the given equation is

"u=u_p+u_{hom}=2e^{-x}+\frac{3}{4}\sin 2y+f_1(x+y)+f_2(x-y)"